OpenStudy (anonymous):

Farmer Bown planted corn and wheat his 1210 acres of land. The cost of planting and arvesting corn (which includes seed, planting, fertilizer, machinery, labor, and other costs) is$275 per acre. The cost of planting wheat is $125 per acre. If Farmer Brown's total cost was $254,750, how many acres of corn did he plant?

OpenStudy (anonymous):

$254,750 = corn*$275 + wheat*$125 corn + what = 1210

OpenStudy (anonymous):

254,750 = corn*275 + wheat*125 1210 = corn + wheat 254,750 = corn*275 + wheat*125 1210*275= corn*275 + wheat*275 now we have a comparison: corn doesn't change, all difference must be wheat

OpenStudy (anonymous):

can also be solved by substitution 254,750 = corn*275 + wheat*125 -- want to get corn alone: 254,750 - wheat*125= corn*275 corn = 254,750/275 - wheat*125/275 now substitute in the other equation: 1210 = corn + wheat 1210 = 254,750/275 - wheat*125/275 + wheat 1210 = 926.36 - wheat*125/275 + wheat 1210 = 926.36 - wheat*125/275 + wheat |(-926.36) 283.64 = -wheat*125/275 + wheat 283.64 = -0.45 wheat + wheat 283.64 = 0.55 wheat |(x100) 28364 = 55 wheat |(:55) wheat = 515 acres